# Consistent Validation for Predictive Methods in Spatial Settings

### v1

@@ B.5.1 Preliminary Lemma: Relating Fill Distances
B.5.2 Bound on Loss Depending on Fill Distance
@@
Corollary B.10 (Bound for Dense Validation Data and General Test Data).
@@
$|R^{\chi}_{Q^{\textup{{test}}}}(h)-\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T_{2}}}}(h)|\leq K_{d,\delta,\Delta,L}\tilde{\rho}^{\frac{d}{d+2}},$ (85)
@@
Choose $k=\min(\lceil(\gamma\tilde{\rho})^{-\frac{2d}{d+2}}\rceil,N^{\textup{{val}}})$ for some $\gamma\in(0,1/2)$ to be specified later. Because $\frac{2d}{d+2}<d$ and $\gamma\leq\frac{1}{2}$, this $k$ satisfies the conditions of Prop. B.9 so,
@@
We now apply Thm. 5.1 and Prop. B.8 together with this bound to conclude with probability $1-\delta$
$|R^{\chi}_{Q^{\textup{{test}}}}(h)-\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T_{2}}}}(h)|$ $\leq\sqrt{2}C_{L}(4\gamma^{-2}\tilde{\rho}^{\frac{d}{d+2}}+\frac{1}{\sqrt{k}}\beta_{\delta})$ (89)
@@
B.5.3 Consistency of Spatial Nearest Neighbors
We now restate and prove that the spatial nearest neighbor procedure we describe is consistent under infill asymptotics. This follows as a corollary of Cor. B.10 since the infill assumption means that for any fixed $\delta$, the upper bound in Cor. B.10 tends to zero with the fill distance.
@@ B.5.3 Consistency of Spatial Nearest Neighbors
Take $\delta=\min(1,\tilde{\rho})$. Under infill asymptotics, this tends to zero. On the other hand, by Cor. B.10 (or more precisely Eqn. 90 which makes the depend of the bound on $\delta$ explicit), with probability at least $1-\delta$
$|R^{\chi}_{Q^{\textup{{test}}}}(h)-\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T_{2}}}}(h)|\leq K_{d,\Delta,L}\tilde{\rho}^{d/(d+2)}\sqrt{\log\frac{1}{\tilde{\rho}}}.$ (91)
The left hand side of this bound tends to zero with $\tilde{\rho}$, and so $\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T_{2}}}}(h)$ converges in probability to $R^{\chi}_{Q^{\textup{{test}}}}(h)$. ∎
B.6 Model Selection: Rates of Convergence of Spatial Nearest Neighbors
We now present an extended version of earlier discussion in Section 5.3, as well as results on rates of convergence of spatial nearest neighbors that provides some support to claims regarding model selection. As a special case, we consider grid prediction (for example for assessing the global performance of a map constructed using a predictive method). We then discuss rates of convergence in the more general setting earlier addressed in Cor. B.10.
B.6.1 A Heuristic Discussion of Model Selection with Increasing Amounts of Data
@@
We will assume we are in the additive, homoskedastic error setting so that $Y=f(S,\chi(S))+\epsilon$. We will assume that the noise is bounded, and that squared loss is used. Finally, we will assume that training data is also generated following this process. We will also assume that $f$ is Lipschitz continuous. In this setting, minimax pointwise regression rates are $\theta((N^{\textup{{train}}})^{-\frac{1}{d+2}})$ (cf. Tsybakov (2008, Theorem 2.3, Corollary 2.2) in one-dimension under a fixed grid design and Tibshirani (2023, Examples 3.1, 3.2) for a multi-dimensional version with both fixed grid and random design. The latter assumes Gaussian noise instead of bounded noise). Ideally, we would like a method for performing model selection to be able to distinguish between a sequence of predictive methods converging at slower than the minimax optimal rate and a sequence of predictive methods converging at the minimax rate and to be able to reliably select between two sequences of predictive methods both of which are converging at the minimax rate, but with different constants.
@@
B.6.2 Rates of Convergence for Grid Prediction
@@
Assumption B.11 (Grid prediction).
@@
Corollary B.12 (Bound on Estimation Error for Grid Prediction).
With the same assumptions as Thm. 5.1 and additionally Assumption B.11, with probability at least $1-\delta$
$\!\!\!|R^{\chi}_{Q^{\textup{{test}}}}(h)\!-\!\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}}(h)|\!\leq\!L\rho\!+\!\beta_{\delta}\sqrt{\max(\tfrac{2^{d}}{M^{\textup{{test}}}},(8\rho)^{d})}.$ (92)
@@
$\!\!|R^{\chi}_{Q^{\textup{{test}}}}(h)\!-\!\!\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T_{2}}}}(h)|\leq\!C_{L}\!\!\left(\!\!\rho\!+\beta_{\delta}\sqrt{\max(\tfrac{2^{d}}{M^{\textup{{test}}}},(8\rho)^{d})}\right).$
@@
See Section B.6.5 for a proof. The right-hand side of the bound is small as long as there is a validation point near each test point and the resolution of the map is high. If the available data for validation is generated i.i.d. as in Prop. 3.3, then right hand side becomes, up to logarithmic factors in $N^{\textup{{val}}}$, $\frac{1}{\sqrt{M^{\textup{{test}}}}}+(N^{\textup{{val}}})^{\min(-1/2,-1/d)}$
B.6.3 Statement and Discussion of Result for IID validation data and Grid Prediction
Corollary B.13 (Convergence of Grid Prediction with Independent and Identically Distributed Validation Data).
Suppose that $\mathcal{S}=[0,1]^{d}$, $S^{\textup{{val}}}_{n}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}P$ for $1\leq n\leq N^{\textup{{val}}}$ with $N^{\textup{{val}}}>1$ and $P$ has Lebesgue density lower bounded by $c>0$. Additionally, take the assumptions of Cor. B.12. Fix $\delta\in(0,1)$ and $k\in\{1,k^{\star}_{T_{2}}\}$. Then there exists a constant $K_{d,\delta,L,\Delta,c}$ that depends only on $d,\delta,L,c$ and $\Delta$ such that with probability at least $1-\delta$
$|\hat{R}^{\chi}_{w^{\textup{{NN}},k}}(h)-R^{\chi}_{Q^{\textup{{test}}}}(h)|\leq K_{d,\delta,L,\Delta,c}\left((\tfrac{\log N^{\textup{{val}}}}{N^{\textup{{val}}}})^{\min(\frac{1}{2},\frac{1}{d})}+\tfrac{1}{\sqrt{M^{\textup{{test}}}}}\right).$
@@ B.6.3 Statement and Discussion of Result for IID validation data and Grid Prediction
From Prop. 3.3, there exists a constant $\gamma$ depending on $d,L,\delta,c,\Delta$ such that with probability at least $21-\delta/$ 2
@@
An upper bound on the bound in Cor. B.12 shows that for $k\in\{1,k^{\star}_{T_{2}}\}$ with probability $21-\delta/$ 2,
$|\hat{R}^{\chi}_{w^{\textup{{NN}},k}}(h)-R^{\chi}_{Q^{\textup{{test}}}}(h)|\leq\gamma C_{L}\max(\beta_{\delta/2}8^{d/2},1)\left(\rho+\frac{1}{\sqrt{M^{\textup{{test}}}}}+\rho^{d/2}\right).$ (94)
Combining Eqn. 93 and Eqn. 94 via a union bound and using that $a+b\leq 2\max(a,b)$ completes the proof. ∎
This matches the bound proven in Portier et al. (2023, Proposition 3) which assumed that the test data was independent and identically distributed instead of on a regular grid. This rate of convergence is reasonably fast, particularly in low-dimensions. In particular ignoring the dependence of the bound on $M^{\textup{{test}}}$, which is reasonable as the number of test points in map prediction is often far larger than the number of available points for training and validation, it is faster than the minimax optimal rate of convergence for Lipschitz functions of $\theta((N^{\textup{{train}}})^{-1/(d+2)})$ discussed earlier. We therefore would expect both spatial nearest neighbors and $1$-nearest neighbors to perform well for model selection for grid prediction tasks if a fixed percentage of the data is used for training, and the remainder for validation. We emphasize we do not give a formal proof of this, just a heuristic argument suggesting why this should be the case. To give a formal proof would involve at least ensuring estimates of the risk estimation procedure hold uniformly over both sequences of predictive methods, and therefore involve additional assumptions.
B.6.4 General Prediction Tasks
For general $Q^{\textup{{test}}}$ we can combine Cor. B.10 together with Prop. 3.3 to get some sense of the rate of convergence of the spatial nearest neighbor method if the validation data is independent and identically distributed form a measure with density lower bounded on $[0,1]^{d}$ and the test task is fixed.
Corollary B.14 (Convergence of Spatial Nearest Neighbor with Independent and Identically Distributed Validation Data).
Suppose that $\mathcal{S}=[0,1]^{d}$, $S^{\textup{{val}}}_{n}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}P$ for $1\leq n\leq N^{\textup{{val}}}$ with $N^{\textup{{val}}}>1$ and $P$ has Lebesgue density lower bounded by $c>0$. Additionally, take the assumptions of Cor. B.10. Fix $\delta\in(0,1)$. Then there exists a constant $K_{d,\delta,L,\Delta,c}$ that depends only on $d,\delta,L,c$ and $\Delta$ such that with probability at least $1-\delta$
$|\hat{R}^{\chi}_{w^{\textup{{NN}},k}}(h)-\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T_{2}}}}(h)|\leq K_{d,\delta,L,\Delta,c}\left((\tfrac{\log N^{\textup{{val}}}}{N^{\textup{{val}}}})\right)^{\frac{1}{d+2}}.$ (95)
@@ B.6.4 General Prediction Tasks
From Prop. 3.3, there exists a constant $\gamma$ depending on $d,L,\delta,c,\Delta$ such that with probability at least $21-\delta/$ 2
@@
Cor. B.10 implies that there exists a constant $K\geq 0$ such that with probability at least $21-\delta/$ 2
$|\hat{R}^{\chi}_{w^{\textup{{NN}},k}}(h)-\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T_{2}}}}(h)|\leq K\tilde{\rho}^{\frac{d}{d+2}}.$ (97)
Combining Eqn. 96 and Eqn. 97 completes the proof. ∎
In this case, again up to logarithmic factors, Cor. B.14 means that SNN converges at the optimal rate of convergence for Lipschitz functions. This means we do not necessarily expect to be able to distinguish between two sequences of predictive methods that converge at the minimax rate, but we might expect to distinguish between two sequences of predictive methods if one converges to the optimal predictor at much slower than the minimax rate. We again emphasize that we do not formally show this, and to do so would would involve at least ensuring estimates of the risk estimation procedure hold uniformly over both sequences of predictive methods, and therefore involve additional assumptions.
@@
B.6.5 Grid Prediction Proofs
In order to prove the claimed upper bound on grid prediction Cor. B.12 we use Thm. 5.1 (in particular Eqn. 13) together with an upper bound on the number of test points that lie within a ball of radius equal to the fill distance around any validation point. In order to do this, we will use that all the points in a grid are well-separated. We therefore begin by recalling the definition of a packing of a set, as well as a relationship between covering number and packing number.
Definition B.15 (Packing, Packing Number).

### v2

@@ D.5.1 Preliminary Lemma: Relating Fill Distances
D.5.2 Bound on Loss Depending on Fill Distance
@@
Corollary D.14 (Bound for Dense Validation Data and General Test Data).
@@
$|R_{Q^{\textup{{test}}}}(h)-\hat{R}_{\textup{{{NN}}},k^{\star}_{T_{2}}}(h)|\leq K_{d,\delta,\Delta,L}\tilde{\rho}^{\frac{d}{d+2}},$ (91)
@@
Choose $k=\min(\lceil(\gamma\tilde{\rho})^{-\frac{2d}{d+2}}\rceil,N^{\textup{{val}}})$ for some $\gamma\in(0,1/2)$ to be specified later. Because $\frac{2d}{d+2}<d$ and $\gamma\leq\frac{1}{2}$, this $k$ satisfies the conditions of Proposition D.13 so,
@@
We now apply Theorem 5.1 and Proposition D.10 together with this bound to conclude with probability $1-\delta$
$|R_{Q^{\textup{{test}}}}(h)-\hat{R}_{\textup{{{NN}}},k^{\star}_{T_{2}}}(h)|$ $\leq\sqrt{2}C_{L}(4\gamma^{-2}\tilde{\rho}^{\frac{d}{d+2}}+\frac{1}{\sqrt{k}}\beta_{\delta})$ (95)
@@
D.5.3 Consistency of Spatial Nearest Neighbors
We now restate and prove that the spatial nearest neighbor procedure we describe is consistent under infill asymptotics. This follows as a corollary of Corollary D.14 since the infill assumption means that for any fixed $\delta$, the upper bound in Corollary D.14 tends to zero with the fill distance.
@@ D.5.3 Consistency of Spatial Nearest Neighbors
Take $\delta=\min(1,r)$. Under infill asymptotics, this tends to zero because $r\leq C\tilde{\rho}$ and $\tilde{\rho}$ tends to 00 by assumption. Corollary D.14 (or more precisely Equation 96 which makes the depend of the bound on $\delta$ explicit), with probability at least $1-\delta$
$|R_{Q^{\textup{{test}}}}(h)-\hat{R}_{\textup{{{NN}}},k^{\star}_{T_{2}}}(h)|$ $\leq K_{d,\Delta,L}\tilde{\rho}^{d/(d+2)}\sqrt{\log\frac{1}{r}}$ (97)
$\leq K_{d,\Delta,L}\tilde{\rho}^{d/(d+2)}\sqrt{\log\frac{1}{c\tilde{\rho}}}.$ (98)
The left hand side of this bound tends to zero with $\tilde{\rho}$, and so $\hat{R}_{\textup{{{NN}}},k^{\star}_{T_{2}}}(h)$ converges in probability to $R_{Q^{\textup{{test}}}}(h)$. ∎
D.5.4 Computation of An Approximate Fill Distance
The fill distance can be computed exactly by computing the vertices of a $d$-dimensional Voronoi diagram with the validation points, then computing the maximum distance from each of these vertices to a point in the validation set. For problems in $1$ or $2$ dimensions, this is feasible even with a large number of validation points, because algorithms for computation of the Voronoi diagram can be done in nearly linear, $O(N^{\textup{{val}}}\log N^{\textup{{val}}})$ time in $1$ and $2$ dimensions [Okabe et al., 2000, Chapter 4]. In higher dimensions, the worst case complexity of algorithms for computing Voronoi diagrams can be $O((N^{\textup{{val}}})^{\lfloor d/2\rfloor}+N^{\textup{{val}}}\log N^{\textup{{val}}})$, and so for large number of points in more the computational cost can become quite high.
To avoid this computational cost, we instead use a simple space partitioning algorithm when $\mathcal{S}=[0,1]^{d}$ that is guaranteed to give an approximation to the fill distance $r$ satisfying $\frac{\tilde{\rho}}{2\sqrt{d}}\leq r\leq 2\tilde{\rho}$. The idea is to split the domain into $2^{d}$ quadrants and check if each quadrant contains a validation point. If it does, recurse, otherwise stop and keep track of the side length of each quadrant, call it $r$. When the algorithm terminates it must be the case that there exists a partitioning of $[0,1]^{d}$ into cubes of side length $2r$, with each cube containing at least one validation point. Because this is a partition, each point in the spatial domain must also be within a cube, and so the fill distance is upper bounded by the maximum distance between two points in a cube of side length $2r$, i.e. $\tilde{\rho}\leq 2r\sqrt{d}$. Also, when the algorithm stops, a cube of side length $r$ has been found that does not contain any points. Therefore the fill distance is lower bounded by the distance from the center of this cube to the closest validation point, which must be at least $r/2$. That is, $\tilde{\rho}\geq r/2$. Rearranging, we see that
$\frac{\tilde{\rho}}{2\sqrt{d}}\leq r\leq 2\tilde{\rho}$ (99)
as claimed. Finally, we address the computational complexity of this approach. The number of times we recurse is $\log_{2}r$, which is $O(\log 1/\tilde{\rho})$, which is in turn $O(\log N^{\textup{{val}}})$, because the fill distance cannot decrease faster than the inverse of the covering number of $[0,1]^{d}$, which certainly does not decrease faster than $1/N^{\textup{{val}}}$.
It remains to consider the complexity of deciding which orthant all of the validation points lie in, and partioning the points by orthant. This can be done by looping over each of the dimensions, partitioning the points in the cube based on whether or not that coordinate is on the left or right hand side of the current cube, which is $O(N^{\textup{{val}}}d)$. Therefore, the total computational complexity of this algorithm is not more than $O(N^{\textup{{val}}}d\log N^{\textup{{val}}})$, which is nearly linear in $N^{\textup{{val}}}$.
D.6 Model Selection: Rates of Convergence of Spatial Nearest Neighbors
We now present an extended version of earlier discussion in Appendix C, as well as results on rates of convergence of spatial nearest neighbors that provides some support to claims regarding model selection. As a special case, we consider grid prediction (for example for assessing the global performance of a map constructed using a predictive method). We then discuss rates of convergence in the more general setting earlier addressed in Corollary D.14.
D.6.1 A Heuristic Discussion of Model Selection with Increasing Amounts of Data
@@
We will assume we are in the additive, homoskedastic error setting so that $Y=f(S,\chi(S))+\epsilon$. We will assume that the noise is bounded, and that squared loss is used. Finally, we will assume that training data is also generated following this process. We will also assume that $f$ is Lipschitz continuous. In this setting, minimax pointwise regression rates are $\theta((N^{\textup{{train}}})^{-\frac{1}{d+2}})$ (cf. Tsybakov [2008, Theorem 2.3, Corollary 2.2] in one-dimension under a fixed grid design and Tibshirani [2023, Examples 3.1, 3.2] for a multi-dimensional version with both fixed grid and random design. The latter assumes Gaussian noise instead of bounded noise). Ideally, we would like a method for performing model selection to be able to distinguish between a sequence of predictive methods converging at slower than the minimax optimal rate and a sequence of predictive methods converging at the minimax rate and to be able to reliably select between two sequences of predictive methods both of which are converging at the minimax rate, but with different constants.
@@
D.6.2 Rates of Convergence for Grid Prediction
@@
Assumption D.15 (Grid prediction).
@@
Corollary D.16 (Bound on Estimation Error for Grid Prediction).
With the same assumptions as Theorem 5.1 and additionally Assumption D.15, with probability at least $1-\delta$
$\!\!\!|R_{Q^{\textup{{test}}}}(h)\!-\!\hat{R}_{\textup{{NN}},1}(h)|\!\leq\!L\rho\!+\!\beta_{\delta}\sqrt{\max(\tfrac{2^{d}}{M^{\textup{{test}}}},(8\rho)^{d})}.$ (100)
@@
$\!\!|R_{Q^{\textup{{test}}}}(h)\!-\!\!\hat{R}_{\textup{{{NN}}},k^{\star}_{T_{2}}}(h)|\leq\!C_{L}\!\!\left(\!\!\rho\!+\beta_{\delta}\sqrt{\max(\tfrac{2^{d}}{M^{\textup{{test}}}},(8\rho)^{d})}\right).$
@@
See Section D.6.5 for a proof. The right-hand side of the bound is small as long as there is a validation point near each test point and the resolution of the map is high. If the available data for validation is generated i.i.d. as in Proposition 3.3, then right hand side becomes, up to logarithmic factors in $N^{\textup{{val}}}$, $\frac{1}{\sqrt{M^{\textup{{test}}}}}+(N^{\textup{{val}}})^{\min(-1/2,-1/d)}$
D.6.3 Statement and Discussion of Result for IID validation data and Grid Prediction
Corollary D.17 (Convergence of Grid Prediction with Independent and Identically Distributed Validation Data).
Suppose that $\mathcal{S}=[0,1]^{d}$, $S^{\textup{{val}}}_{n}\stackrel{{\scriptstyle\textup{{\tiny iid}}}}{{\sim}}P$ for $1\leq n\leq N^{\textup{{val}}}$ with $N^{\textup{{val}}}>1$ and $P$ has Lebesgue density lower bounded by $c>0$. Additionally, take the assumptions of Corollary D.16. Fix $\delta\in(0,1)$ and $k\in\{1,k^{\star}_{T_{2}}\}$. Then there exists a constant $K_{d,\delta,L,\Delta,c}$ that depends only on $d,\delta,L,c$ and $\Delta$ such that with probability at least $1-\delta$
$|\hat{R}_{\textup{{NN}},k}(h)-R_{Q^{\textup{{test}}}}(h)|\leq K_{d,\delta,L,\Delta,c}\left((\tfrac{\log N^{\textup{{val}}}}{N^{\textup{{val}}}})^{\min(\frac{1}{2},\frac{1}{d})}+\tfrac{1}{\sqrt{M^{\textup{{test}}}}}\right).$
@@ D.6.3 Statement and Discussion of Result for IID validation data and Grid Prediction
From Proposition 3.3, there exists a constant $\gamma$ depending on $d,L,\delta,c,\Delta$ such that with probability at least $21-\delta/$ 2
@@
An upper bound on the bound in Corollary D.16 shows that for $k\in\{1,k^{\star}_{T_{2}}\}$ with probability $21-\delta/$ 2,
$|\hat{R}_{\textup{{NN}},k}(h)-R_{Q^{\textup{{test}}}}(h)|\leq\gamma C_{L}\max(\beta_{\delta/2}8^{d/2},1)\left(\rho+\frac{1}{\sqrt{M^{\textup{{test}}}}}+\rho^{d/2}\right).$ (102)
Combining Equation 101 and Equation 102 via a union bound and using that $a+b\leq 2\max(a,b)$ completes the proof. ∎
This matches the bound proven in Portier et al. [2023, Proposition 3] which assumed that the test data was independent and identically distributed instead of on a regular grid. This rate of convergence is reasonably fast, particularly in low-dimensions. In particular ignoring the dependence of the bound on $M^{\textup{{test}}}$, which is reasonable as the number of test points in map prediction is often far larger than the number of available points for training and validation, it is faster than the minimax optimal rate of convergence for Lipschitz functions of $\theta((N^{\textup{{train}}})^{-1/(d+2)})$ discussed earlier. We therefore would expect both spatial nearest neighbors and $1$-nearest neighbors to perform well for model selection for grid prediction tasks if a fixed percentage of the data is used for training, and the remainder for validation. We emphasize we do not give a formal proof of this, just a heuristic argument suggesting why this should be the case. To give a formal proof would involve at least ensuring estimates of the risk estimation procedure hold uniformly over both sequences of predictive methods, and therefore involve additional assumptions.
D.6.4 General Prediction Tasks
For general $Q^{\textup{{test}}}$ we can combine Corollary D.14 together with Proposition 3.3 to get some sense of the rate of convergence of the spatial nearest neighbor method if the validation data is independent and identically distributed form a measure with density lower bounded on $[0,1]^{d}$ and the test task is fixed.
Corollary D.18 (Convergence of Spatial Nearest Neighbor with Independent and Identically Distributed Validation Data).
Suppose that $\mathcal{S}=[0,1]^{d}$, $S^{\textup{{val}}}_{n}\stackrel{{\scriptstyle\textup{{\tiny iid}}}}{{\sim}}P$ for $1\leq n\leq N^{\textup{{val}}}$ with $N^{\textup{{val}}}>1$ and $P$ has Lebesgue density lower bounded by $c>0$. Additionally, take the assumptions of Corollary D.14. Fix $\delta\in(0,1)$. Then there exists a constant $K_{d,\delta,L,\Delta,c}$ that depends only on $d,\delta,L,c$ and $\Delta$ such that with probability at least $1-\delta$
$|\hat{R}_{\textup{{NN}},k}(h)-\hat{R}_{\textup{{{NN}}},k^{\star}_{T_{2}}}(h)|\leq K_{d,\delta,L,\Delta,c}\left((\tfrac{\log N^{\textup{{val}}}}{N^{\textup{{val}}}})\right)^{\frac{1}{d+2}}.$ (103)
@@ D.6.4 General Prediction Tasks
From Proposition 3.3, there exists a constant $\gamma$ depending on $d,L,\delta,c,\Delta$ such that with probability at least $21-\delta/$ 2
@@
Corollary D.14 implies that there exists a constant $K\geq 0$ such that with probability at least $21-\delta/$ 2
$|\hat{R}_{\textup{{NN}},k}(h)-\hat{R}_{\textup{{{NN}}},k^{\star}_{T_{2}}}(h)|\leq K\tilde{\rho}^{\frac{d}{d+2}}.$ (105)
Combining Equation 104 and Equation 105 completes the proof. ∎
In this case, again up to logarithmic factors, Corollary D.18 means that SNN converges at the optimal rate of convergence for Lipschitz functions. This means we do not necessarily expect to be able to distinguish between two sequences of predictive methods that converge at the minimax rate, but we might expect to distinguish between two sequences of predictive methods if one converges to the optimal predictor at much slower than the minimax rate. We again emphasize that we do not formally show this, and to do so would would involve at least ensuring estimates of the risk estimation procedure hold uniformly over both sequences of predictive methods, and therefore involve additional assumptions.
@@
D.6.5 Grid Prediction Proofs
In order to prove the claimed upper bound on grid prediction Corollary D.16 we use Theorem 5.1 together with an upper bound on the number of test points that lie within a ball of radius equal to the fill distance around any validation point. In order to do this, we will use that all the points in a grid are well-separated. We therefore begin by recalling the definition of a packing of a set, as well as a relationship between covering number and packing number.
Definition D.19 (Packing, Packing Number).